# TGB 2.0: A Benchmark for Learning on Temporal Knowledge Graphs and Heterogeneous Graphs

### v1

@@ Appendix B Broader Impact
Impact on Temporal Graph Learning. Recently, the availability of large graph benchmarks accelerates research in the field [25, 24, 10]. By providing a standardized benchmarking framework, TGB 2.0 will accelerate the development and evaluation of new models for temporal knowledge graphs and temporal heterogeneous graphs. Researchers can build upon a common foundation, leading to more rapid and robust advancements in this field. In addition, the introduction of a unified evaluation framework addresses reproducibility issues, which are critical for scientific progress. The comprehensive evaluation facilitated by TGB 2.0 ensures that new methods are rigorously tested against state-of-the-art baselines, leading to more robust and well-validated models. This contributes to higher standards in research and more reliable outcomes. Overall, this work has the potential to significantly impact both the academic research community and practical applications, driving forward the understanding and utilization of multi-relational temporal graphs in various fields.
@@
Potential Negative Impact. The TGB 2.0 datasets may limit the utilization and mining of other TG datasets. If the datasets are not representative of the broader set of real-world data, this could lead to biased or unfair outcomes when models are applied in practice. Similarly, the community might become overly dependent on the TGB 2.0 framework, potentially hindering the exploration of alternative benchmarking methodologies or the development of diverse evaluation protocols that might be more suitable for specific contexts or emerging subfields. Moreover, when the focus is mainly on quantitative performance metrics, it might overshadow the importance of qualitative assessments and other critical factors such as interpretability, fairness, and ethical considerations in model development and deployment. To avoid this issue, we plan to update TGB regularly with community feedback as well as adding additional datasets and tasks.
Appendix C Dataset Documentation and Intended Use
@@
All datasets presented by TGB 2.0 are intended for academic use and their corresponding licenses are listed in Appendix D. We also anonymized the datasets, to remove any personally identifiable information where appropriate. For the ease of access, we provide the following links to the TGB 2.0 benchmark suits and datasets.
@@
Appendix D Dataset Licenses and Download Links
@@ Appendix E Dataset Statistics
Figure 7 shows the top ten most frequent edge types in TKG datasets. Figure 8 shows the top ten most frequent edge types in THG datasets. Note that TKG datasets in general has more edge types than THG datasets. Most common THG relations usually share similar portion of edges in the dataset while TKG relations shares different portion of edges.
@@ F.1 Computing Resources
We ran all experiments on either Narval or Béluga cluster of Digital Research Alliance of Canada or the Mila, Québec AI Institute cluster. For the experiments on the Narval cluster, we ran each experiment on a Nvidia A100 (40G memory) GPU with 4 CPU nodes (from either of the AMD Rome 7532 @ 2.40 GHz 256M cache L3, AMD Rome 7502 @ 2.50 GHz 128M cache L3, or AMD Milan 7413 @ 2.65 GHz 128M cache L3 available type) each with 100GB memory. For experiments on the Béluga cluster, we rand each experiments on a NVidia V100SXM2 (16G memory) GPU wiht 4 CPU nodes (from Intel Gold 6148 Skylake @ 2.4 GHz) each with 100GB memory. For the experiments on the Mila cluster, we ran each experiment on an RTX8000 (40G memory) GPU or an V100 (32G memory) GPU with 4 CPU nodes (from either of the AMD Rome 7532 @ 2.40 GHz 256M cache L3, AMD Rome 7502 @ 2.50 GHz 128M cache L3, or AMD Milan 7413 @ 2.65 GHz 128M cache L3 available type). The upper limit of RAM was set to 1056GB.
A seven-day time limit was considered for each experiment. For all non deterministic methods, i.e. all methods besides Edgebank and the Recurrency Baseline, we repeated each experiments five times and reported the average and standard deviation of different runs. It is noteworthy that except for the reported baseline results, the other models, all evaluated by their original source code, throw an out of memory error or do not finish in the given time limit for the medium and large datasets on all available resources includiing Narval, Béluga, and Mila clusters.
@@ F.3.1 Hyperparameters
If not stated otherwise, for each method we use the hyperparameter setting as reported in the original papers, please see Table 8. Whereas further hyperparameter tuning could further improve performance of each method, it was out of scope for this work. We only change the hyperparameter values only if the methods would not finish with the given time or memory limit. In this case, we follow recommendations from [14] (to decrease rule length and window size for TLogic), from the authors of [13] (to decrease the window length for the Recurrency Baseline), and from the authors of [39] (to decrease the history length for RE-GCN and CEN).
Table 8: Hyperparameter choices. Values that are different from the original papers are bolded. In case we modify the values for different datasets, we report so in the respective columns.
@@
RecB $\lambda=0.1$, $\alpha=0.99$, $\text{window}=0$ tkgl-icews: $\text{window}=500$
All Datasets Dataset-specific
RecB $\lambda=0.1$, $\alpha=0.99$, $\text{window}=0$
@@ F.5 Ablation Study on Negative Sample Generation
1-vs-all $\text{RecB}_{\text{default}}$ [13] 0.640 0.570 316 659
@@
F.6 Detailed information on Train, Validation, and Test Splits
As described in Section 4, we split all datasets chronologically into the training, validation, and test sets, respectively containing $70\%$, $15\%$, and $15\%$ of all edges. Because we ensure that edges for a timesteps can only be in either train or validation or test set, and because the number of edges over time are not constant, the cuts are not strict. We provide more details on the exact splits in Table 10.
@@ G.2 Temporal Heterogeneous Graph Forecasting
  • TGN [58] represents a comprehensive framework designed for learning on dynamic graphs in continuous time. Its components include a memory module, message function, message aggregator, memory updater, and embedding module. During testing, TGN updates the memories of nodes with edges that have been newly observed. Additionally, to incorporate edge types into the TGN, we devised a variant of the TGN capable of utilizing edge type information. This was achieved by generating embeddings from the edge types, which were then concatenated with the original messages within the TGN model.
@@ H.0.3 Collection Process
    The data is extracted from online public data sources. The data described different relations between entities. The data sources are found in Appendix D and dataset details are in Section 4.
@@
    For thgl-myket, the users where selected randomly among the users that have interactions with the platform in a two-week period. For tkgl-smallpedia, tkgl-wikidata, the dataset was filtered by Wiki page ID. thgl-software and thgl-github, nodes with low degrees are filtered out.
@@
    All datasets are obtained via websites except for thgl-myket which were provided by the the Myket Android application market team. Links to data sources are in Appendix D.
@@ H.0.5 Uses
    Yes, all paper references are provided in this paper. All data sources are discussed in Appendix D.
@@
    No and we discuss potential negative impacts in Appendix B.
@@ H.0.6 Distribution
    Yes, the DOI for the project is https://zenodo.org/records/11480522 (will point to all future version as well). The dataset download links are provided in Appendix D. TGB 2.0 datasets are maintained via Digital Research Alliance of Canada (funded by the Government of Canada).
@@
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions. The dataset licenses are listed in Appendix D.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions. All license terms are discussed in Appendix D.

### v2

@@ Appendix A Broader Impact
Impact on Temporal Graph Learning. Recently, the availability of large graph benchmarks accelerates research in the field [27, 26, 12]. By providing a standardized benchmarking framework, TGB 2.0 will accelerate the development and evaluation of new models for temporal knowledge graphs and temporal heterogeneous graphs. Researchers can build on a common foundation, leading to more rapid and robust advancements in this field. In addition, the introduction of a unified evaluation framework addresses reproducibility issues, which are critical to scientific progress. The comprehensive evaluation facilitated by TGB 2.0 ensures that new methods are rigorously tested against state-of-the-art baselines, leading to more robust and well-validated models. This contributes to higher standards in research and more reliable outcomes. Overall, this work has the potential to significantly impact both the academic research community and practical applications, driving forward the understanding and utilization of multi-relational temporal graphs in various fields.
@@
Potential Negative Impact. The TGB 2.0 datasets may limit the utilization and mining of other TG datasets. If the datasets are not representative of the broader set of real-world data, this could lead to biased or unfair outcomes when models are applied in practice. Similarly, the community might become overly dependent on the TGB 2.0 framework, potentially hindering the exploration of alternative benchmarking methodologies or the development of diverse evaluation protocols that might be more suitable for specific contexts or emerging sub-fields. In addition, when the focus is mainly on quantitative performance metrics, it could overshadow the importance of qualitative assessments and other critical factors, such as interpretability, fairness, and ethical considerations in the development and implementation of models. To avoid this issue, we plan to update TGB 2.0 regularly with community feedback as well as adding additional datasets and tasks.
Appendix B Dataset Documentation and Intended Use
@@
All datasets presented by TGB 2.0 are intended for academic use and their corresponding licenses are listed in Appendix C. We also anonymized the datasets, to remove any personally identifiable information where appropriate. For ease of access, we provide the following links to the TGB 2.0 benchmark suits and datasets.
@@
Appendix C Dataset Licenses and Download Links
@@
Appendix D Data Processing Details
Here we discuss the data collection and cleaning process for TGB 2.0 datasets in detail. For reproducibility, we also provide the corresponding mining script on Github.
TKG Dataset Format. For all TKG datasets, we include the temporal links in edgelist.csv. The validation and test negative samples are included in val_ns.pkl and test_ns.pkl files respectively. Additionally, for tkgl-smallpedia and tkgl-wikidata, we also include the static links related to the wiki entities in static_edgelist.csv.
THG Dataset Format. For all THG datasets, we include the temporal links in edgelist.csv. The validation and test negative samples are included in val_ns.pkl and test_ns.pkl files respectively. We also include nodemapping.csv and edgemapping.csv to provide the name of each node relation and edge relation respectively. Lastly, nodetype.csv provides the description for node type information.
tkgl-wikidata (scripts). This TKG dataset is extracted from the Wikidata KG [68] from 2024 February 20th. We extract the relations from the first 32 million Wikidata entities (by the entity ID). We then retain the relations with the temporal qualifier with relation IDs: P585, P580, P582, P577, P574. We also retain the static relations with these wiki entities, filtering out less informative relations including P31 and P279. We then keep only links with both start and end date as well as point in time relations. We retain links starting from 0 BC. We also remove any links that show up as both static and temporal link.
tkgl-smallpedia (scripts). This is a subset of the tkgl-wikidata where links from the first 1 million entities are retained and then filtered by their year. Only links from 1900 to 2024 are kept. We also extract the static relations associated with these first 1 million entities. Further, we remove any links that show up as both static and temporal link.
tkgl-polecat (scripts). We extract the raw files from the POLECAT dataset source [62], we organize the monthly edgelist chronologically and then merge them into a single file. Then, we remove any links with missing source or destination.
tkgl-icews (scripts). We extract the raw files from ICEWS Coded Event Data source [7, 65]. We first combine the monthly files into a single one, removing any links missing a source or destination.
thgl-github (scripts). The raw data is extracted from GH Arxiv, containing the data from March 2024. We then extract 14 relations based on the common activities on Github. We also removed the issue comment and pr review comment node types as they are often one time nodes which rarely repeat, and kept nodes which have at least two edges in the dataset.
thgl-software (scripts). The raw data is extracted from GH Arxiv, containing the data from January 2024. We kept nodes with at least 10 interactions while removing the issue comment and pr review comment node types.
thgl-forum (scripts). The raw data is downloaded from here [54]. We first merge edge files and attribute files by edge ID. Next, we filter out nodes with less than 100 links from the dataset. The node types are user nodes and subreddit nodes. The edge types are user replying to user and user posting on subreddit.
thgl-myket (scripts). This is an original dataset provided by the Myket Corporation. The data indicates if an edge is a software upgrade or a first time install which forms the two edge types in this data. The node types are users and apps.
@@ Appendix E Dataset Statistics
Figure 7 shows the top ten most frequent edge types in TKG datasets. Figure 8 shows the top ten most frequent edge types in THG datasets. Note that TKG datasets in general have more edge types than THG datasets. Most common THG relations usually share similar portion of edges in the dataset while TKG relations shares different portion of edges.
@@ F.1 Computing Resources
We ran all experiments on either Narval or Béluga cluster of Digital Research Alliance of Canada or the Mila, Québec AI Institute cluster. For the experiments on the Narval cluster, we ran each experiment on a Nvidia A100 (40G memory) GPU with 4 CPU nodes (from either of the AMD Rome 7532 @ 2.40 GHz 256M cache L3, AMD Rome 7502 @ 2.50 GHz 128M cache L3, or AMD Milan 7413 @ 2.65 GHz 128M cache L3 available type) each with 100GB memory. For experiments on the Béluga cluster, we ran each experiments on a NVidia V100SXM2 (16G memory) GPU wiht 4 CPU nodes (from Intel Gold 6148 Skylake @ 2.4 GHz) each with 100GB memory. For the experiments on the Mila cluster, we ran each experiment on an RTX8000 (40G memory) GPU or an V100 (32G memory) GPU with 4 CPU nodes (from either of the AMD Rome 7532 @ 2.40 GHz 256M cache L3, AMD Rome 7502 @ 2.50 GHz 128M cache L3, or AMD Milan 7413 @ 2.65 GHz 128M cache L3 available type). The upper limit of RAM was set to 1056GB.
A seven-day time limit was considered for each experiment. For all non deterministic methods, i.e. all methods besides Edgebank and the Recurrency Baseline, we repeated each experiments five times and reported the average and standard deviation of different runs. It is noteworthy that except for the reported baseline results, the other models, all evaluated by their original source code, throw an out of memory error or do not finish in the given time limit for the medium and large datasets on all available resources including Narval, Béluga, and Mila clusters.
@@ F.3.1 Hyperparameters
If not stated otherwise, for each method we use the hyperparameter setting as reported in the original papers, please see Table 8. Whereas further hyperparameter tuning could further improve performance of each method, it was out of scope for this work. We only change the hyperparameter values if the methods would not finish with the given time or memory limit. In this case, we follow recommendations from [16] (to decrease rule length and window size for TLogic), from the authors of [15] (to decrease the window length for the Recurrency Baseline), and from the authors of [41] (to decrease the history length for RE-GCN and CEN).
Table 8: Hyperparameter choices for each method for All datasets. Values that are different from the original papers are bolded. In case we modify the values for different datasets, we report so in the column Dataset-specific.
@@
$\text{RecB}_{\text{default}}$ $\lambda=0.1$, $\alpha=0.99$, $\text{window}=0$ tkgl-icews: $\text{window}=500$
[…]
@@ F.5 Ablation Study on Negative Sample Generation
1-vs-all $\text{RecB}_{\text{default}}$ [15] 0.542 0.486 316 659
@@
F.6 Details on Evaluation Protocol
As described in section 1, we compute the time-aware Mean Reciprocal Rank [16]. Specifically, for each test edge $(s_{test},r_{test},o_{test},t_{test})$, we evaluate the models prediction by removing the object $o_{test}$ from the test query, $(s_{test},r_{test},?,t_{test})$. The model then assigns scores to all possible entities for the object position, which are subsequently ranked in descending order. This is repeated for subject prediction. The MRR is calculated as the mean of the reciprocal of these ranks across all test queries. In applying the time-aware filter setting, we filter out quadruples that have the same timestamp as the test query. For example, for a test query (France, wins, Basketball Match, 2025-05-19) a model’s prediction (France, wins, Soccer Match, 2025-05-19) would be excluded if (France, wins, Soccer Match, 2025-05-19) is present in the test set. However, it would not be filtered out if (France, wins, Soccer Match, 2025-03-09) is present in the test set.
We evaluate all methods in single-step prediction. This means that the model always forecasts the next timestep, and then ground truth facts are fed before predicting the subsequent timestep [16]. If there are multiple triples in the candidate set with the same score from the model, our protocol assigns the average rank, i.e. the average between the worst (pessimistic) and best (optimistic) score, to the ground truth triple.
The scripts for generating the negative samples can be found here under the folder for each individual dataset 333the scripts all have suffix of ns_gen.py. We select the following negative sampling strategies for each dataset:
  • tkgl-smallpedia, tkgl-polecat, tkgl-icews: 1-vs-all
  • tkgl-wikidata, thgl-software: 1-vs-q with $q=1000$
  • thgl-github, thgl-myket: 1-vs-q with $q=20$
  • thgl-forum: 1-vs-q with $q=100$
F.7 Results on Additional Metrics
In addition to the MRR, in Table 10, we report the Hits at 10 (H@10), the proportion of test queries for which at least one correct item is among the top 10 ranked results. The results show that the ranking of methods in Hits@10 and MRR metric are closely, but not fully matched.
Table 10: Comparison of test results when using Hits at 10 (H@10) vs. MRR metric for Temporal Knowledge Graph Link Prediction task.
Method tkgl-smallpedia tkgl-polecat tkgl-icews tkgl-wikidata
Test Score MRR H@10 MRR H@10 MRR H@10 MRR H@10
$\text{EdgeBank}_{\text{tw}}$ [57] 0.353 0.566 0.056 0.119 0.020 0.058 0.535 0.596
$\text{EdgeBank}_{\infty}$ [57] 0.333 0.562 0.045 0.094 0.009 0.014 0.535 0.596
[…]
RE-GCN [41] 0.594 0.687 0.175 0.292 0.182 0.331 - -
CEN [39] 0.612 0.705 0.184 0.323 0.187 0.334 - -
TLogic [47] 0.595 0.707 0.228 0.378 0.186 0.301 - -
F.8 Detailed information on Train, Validation, and Test Splits
As described in Section 4, we split all datasets chronologically into the training, validation, and test sets, respectively containing $70\%$, $15\%$, and $15\%$ of all edges. Because we ensure that edges for a timestep can only be in either train or validation or test set, and because the number of edges over time are not constant, the cuts are not strict. We provide more details on the exact splits in Table 11.
@@ G.2 Temporal Heterogeneous Graph Forecasting
  • TGN [61] represents a comprehensive framework designed for learning on dynamic graphs in continuous time. Its components include a memory module, message function, message aggregator, memory updater, and embedding module. During testing, TGN updates the memories of nodes with edges that have been newly observed. Additionally, to incorporate edge types into TGN, we devised a variant of TGN capable of utilizing edge type information. This was achieved by generating embeddings from the edge types, which were then concatenated with the original messages within the TGN model.
@@ H.0.3 Collection Process
    The data is extracted from online public data sources. The data described different relations between entities. The data sources are found in Appendix C and dataset details are in Section 4.
@@
    For thgl-myket, the users where selected randomly among the users that have interactions with the platform in a two-week period. For tkgl-smallpedia, tkgl-wikidata, the dataset was filtered by Wiki page ID. For thgl-software and thgl-github, nodes with low degrees are filtered out.
@@
    All datasets are obtained via websites except for thgl-myket which were provided by the the Myket Android application market team. Links to data sources are in Appendix C.
@@ H.0.5 Uses
    Yes, all paper references are provided in this paper. All data sources are discussed in Appendix C.
@@
    No and we discuss potential negative impacts in Appendix A.
@@ H.0.6 Distribution
    Yes, the DOI for the project is https://zenodo.org/records/11480522 (will point to all future version as well). The dataset download links are provided in Appendix C. TGB 2.0 datasets are maintained via Digital Research Alliance of Canada (funded by the Government of Canada).
@@
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions. The dataset licenses are listed in Appendix C.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions. All license terms are discussed in Appendix C.